Simplify 3/(2x-4) + 2/(6-3x)
Answer by palanisamy(496)

3/(2x-4)+2/(6-3x)=3/2(x-2)-2/3(x-2)
=[3*3-2*2]/3*2*(x-2)
=5/6(x-2)